OpenStudy (anonymous):

32 is 80% of what number? - will medal A.0.4 B.4 C.25.6 D.40

OpenStudy (anonymous):

d. 40

OpenStudy (anonymous):

D. 40

OpenStudy (anonymous):

Is correct

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

Divide 32 by 4 = 8 so 32 + 8 = 40

OpenStudy (anonymous):

let the no.=x 80 % of x=3 \[\frac{ 80 }{ 100 }*x=32,x=32*\frac{ 100 }{ 80 }=?\]

OpenStudy (anonymous):

it is not 3,but 32